OpenStudy (anonymous):

How do i find the inverse? I need to find the inverse of f(x) = 2x- 8? How do i do it.

OpenStudy (anonymous):

like solving \[2x-8=20\] add 8 divide by 2 so inverse is \[f^{-1}(x)=\frac{x+8}{2}\]

OpenStudy (anonymous):

Where did you get the 20 from

OpenStudy (amistre64):

a common strategy is to swap the places of x and f(x) and rearrange the setup to solve for f(x)

OpenStudy (anonymous):

is that the right answer though

OpenStudy (amistre64):

if you want something rigorous, you might wanna google it and pull it off one of those .edu sites

OpenStudy (amistre64):

intuitively; the inverse swaps y=x positions

OpenStudy (anonymous):

i made up the 20, idea is that the inverse solves for the variable when you know the output. so method is "add 8, divide by 2" you can always write \[2x-8=y\] \[2x=y+8\] \[x=\frac{y+8}{2}\] so inverse is \[f^{-1}(x)=\frac{x+8}{2}\] or as amistre said you can swap x and y and write \[2y-8=x\] \[2y=x+8\] \[y=\frac{x+8}{2}\] same thing

OpenStudy (anonymous):

thanks